OpenStudy (anonymous):

Solve for x: (x - 9)2 = 1

OpenStudy (anonymous):

x: (x - 9)^2 = 1 Is your question this one

OpenStudy (mertsj):

Why is it that you don't answer the people who try to help you?

OpenStudy (anonymous):

sorry yes that is I read it wrong

OpenStudy (anonymous):

x: (x - 9)^2 = 1 so this is question

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

(x - 9)^2 = 1 (x - 9) (x - 9) = 1 Now expand the brackets. (x - 9) (x - 9) = 1 x^2 - 18x + 81 = 1 Move the 1 across the equal sign x^2 - 18x + 81 = 1 x^2 - 18x + 80 = 0 Factor x^2 - 18x + 80 = 0 (x - 8) (x - 10) = 0 x = 8 or 10

OpenStudy (anonymous):

i hope this answer satisfies you

OpenStudy (anonymous):

method 2: \[(x-9)^2=1\\ \therefore x-9=\pm1\\ \therefore x-9=-1\qquad\qquad x-9=1\\ \therefore x=8\qquad\qquad\qquad x=10 \]